Statistics for Fundamental Algorithms Course

The statistics for the homeworks and the midterm is as follows:

Average Median

HW1  42.64  48
HW2  35.1  41.5
MIDTERM  37.98  38

The distribution for the midterm is as follows:

Range # students 0 − 19 7 20 − 29 8 30 − 39 19 40 − 49 18 50 − 59 11 60 − 69 3 70 − 79 3

In total, 69 students appeared for the midterm exam.